A sample of krypton at a temperature of 125.0 oC occupies 3570. mL. At what temperature in oC will the gas occupy 9.800 L?
Answer #1

Assume Krypton is an ideal gas:

PV1 = nRT1 ...V1=3570mL

PV2 = nRT2 ...V2=9.8*10^3 mL

Devide first equation by second

V1/V2 = T1/T2 ...because P, n and R are constant.

T2 = T1*(V2/V1) = (125+273)*((9.8*10^3)/3570) = 1092.54 K = 819.54 oC.

Thus at 819.5oC, the gas will have 9.800 L volume.